OpenStudy (esshotwired):

Medal for correct answers. 2+2= 3+3= 7+7= Hint: the answer is not what you think it is.

OpenStudy (esshotwired):

fyi: this is just for fun and free medal

OpenStudy (anonymous):

4

OpenStudy (anonymous):

6

OpenStudy (anonymous):

14

OpenStudy (esshotwired):

no i just said, it is not what you think it is. meaning its not that easy

OpenStudy (anonymous):

overall 24

OpenStudy (anonymous):

So your sayin its not 2+2=4 3+3=6 7+7=14 ?

OpenStudy (esshotwired):

yes

OpenStudy (schrodingers_cat):

How about this lol 2 backwards plus 2 = like a fish? 3 backwards plus 3 = 8 7 backwards plus 7 = upside down traingle

OpenStudy (esshotwired):

correcto!

OpenStudy (schrodingers_cat):

lol

OpenStudy (esshotwired):

i do this at least once a week to be nice and see how many people get it

OpenStudy (anonymous):

... Smart man... ;)

OpenStudy (anonymous):

or fish,8, triangle